# Designing AI-Agents with Personalities: A Psychometric Approach

### v3

@@ 3.1.2 Data Analysis
Finally, we conducted Confirmatory Factor Analyses (CFA)  [^4] CFA is a widely used psychometric method for testing whether a set of questionnaire items reflects an expected structure of psychological traits[89]. For example, in the Big Five model, items about being “outgoing” or “sociable” should cluster together as indicators of Extraversion. CFA formally tests such hypotheses by estimating how strongly each item loads onto a latent factor (e.g., Extraversion) and by providing indices of overall model fit (e.g., CFI, RMSEA, SRMR). Good fit means the data are consistent with the hypothesized trait structure, while poor fit suggests the responses may not capture the intended personality dimensions. In our study, CFA allows us to evaluate whether AI-Agent responses reproduce the same underlying Big Five structure observed in human data.using the lavaan package in R [106]. For each LLM, prompt condition, and personality trait, we fitted a one-factor model and compute the one-factor reliability coefficient (a.k.a., Omega). To evaluate the fit of the two CFA models, we used the chi-square test of fit and three approximate fit indices with common cutoff points: 1) the comparative fit index (CFI), with a value above .90 indicating a reasonable fit and above .95 indicating a very good fit; 2) the root mean square error of approximation (RMSEA), with a value of less than .08 indicating reasonable fit and less than .05 indicating very good fit; and 3) the standardized root mean square residual (SRMR), with a value of less than .08 indicating reasonable fit and less than .05 indicating very good fit.
@@ 3.2.1 Methods and Data Analyses
We generated synthetic BFI-2 data based on sample statistics obtained from [112] empirical sample. To do so, we first extracted key statistics from each personality domain. Specifically, within each domain, we extracted (1) the means and standard deviations for each of the three facets, (2) the $3\times 3$ correlation matrix among the three facets, and (3) average correlation among the four items within each facet (See Supplementary Materials for Sample Code).
@@ 4.1.1 Measurement
The five moral dilemma vignettes were designed to assess individuals’ empathetic tendency versus rule adherence tendency. Given that the majority of currently available LLMs have been safety-aligned and instruction fine-tuned [72], which typically skews them away from engaging in severe moral judgments (e.g., the Trolley Problem), we introduced a series of everyday moral dilemmas. These dilemmas necessitated choosing between upholding a standard and prioritizing empathy. Examples include decisions on 1) reporting a friend for cheating on a quiz, 2) addressing a colleague’s misappropriation of office supplies, 3) underage drinking, 4) disclosing confidential information that could save lives, and 5) providing candid feedback on subpar performance (see Appendix III). For the moral dilemmas vignettes, higher scores indicate higher rule adherence tendency and lower empathetic tendency.
@@ 5.1 Limitations and Future Directions
There are several limitations to our study. One limitation is that our method of simulating AI-Agents using only prompts designed from the BFI-2 falls short of capturing the intricate interplay of other important personality constructs, as well as individual backgrounds and societal complexity. We acknowledge that personality traits, while central to much of personality psychology, represent only one facet of what most people consider to be “personality.” Traits capture relatively stable patterns of thinking, feeling, and behaving, but they do not fully encompass other key components such as values, goals, roles, social identity, and emotional dynamics. Recent work has begun to investigate these other aspects in LLMs, such as how LLMs represent emotions [91], human values [122], or identity [116]. These studies highlight that personality in artificial agents can, in principle, extend far beyond trait-level description. Building on these studies, a more integrative approach to LLM personality simulation could combine trait-based control with frameworks for modeling values, motivational systems, and identity-related processes. Such an approach would allow researchers to capture not only how an agent behaves across contexts, but also why it behaves that way, and how it might adapt in response to changing goals or social roles. Expanding in this direction could make AI-Agents a richer platform for studying complex aspects of human personality, bridging dispositional, motivational, and narrative analyses.

### v4

@@ 3.1.2 Data Analysis
Finally, we conducted Confirmatory Factor Analyses (CFA)  [^4] CFA is a widely used psychometric method for testing whether a set of questionnaire items reflects an expected structure of psychological traits[kline2014easy]. For example, in the Big Five model, items about being “outgoing” or “sociable” should cluster together as indicators of Extraversion. CFA formally tests such hypotheses by estimating how strongly each item loads onto a latent factor (e.g., Extraversion) and by providing indices of overall model fit (e.g., CFI, RMSEA, SRMR). Good fit means the data are consistent with the hypothesized trait structure, while poor fit suggests the responses may not capture the intended personality dimensions. In our study, CFA allows us to evaluate whether AI-Agent responses reproduce the same underlying Big Five structure observed in human data.using the lavaan package in R [rosseel2012lavaan]. For each LLM, prompt condition, and personality trait, we fitted a one-factor model and compute the one-factor reliability coefficient (a.k.a., Omega). To evaluate the fit of the two CFA models, we used the chi-square test of fit and three approximate fit indices with common cutoff points: 1) the comparative fit index (CFI), with a value above .90 indicating a reasonable fit and above .95 indicating a very good fit; 2) the root mean square error of approximation (RMSEA), with a value of less than .08 indicating reasonable fit and less than .05 indicating very good fit; and 3) the standardized root mean square residual (SRMR), with a value of less than .08 indicating reasonable fit and less than .05 indicating very good fit.
@@ 3.2.1 Methods and Data Analyses
We generated synthetic BFI-2 data based on sample statistics obtained from [soto_next_2017] empirical sample. To do so, we first extracted key statistics from each personality domain. Specifically, within each domain, we extracted (1) the means and standard deviations for each of the three facets, (2) the $3\times 3$ correlation matrix among the three facets, and (3) average correlation among the four items within each facet (See Supplementary Materials for Sample Code).
@@ 4.1.1 Measurement
The five moral dilemma vignettes were designed to assess individuals’ empathetic tendency versus rule adherence tendency. Given that the majority of currently available LLMs have been safety-aligned and instruction fine-tuned [biedma2024beyond], which typically skews them away from engaging in severe moral judgments (e.g., the Trolley Problem), we introduced a series of everyday moral dilemmas. These dilemmas necessitated choosing between upholding a standard and prioritizing empathy. Examples include decisions on 1) reporting a friend for cheating on a quiz, 2) addressing a colleague’s misappropriation of office supplies, 3) underage drinking, 4) disclosing confidential information that could save lives, and 5) providing candid feedback on subpar performance (see Appendix III). For the moral dilemmas vignettes, higher scores indicate higher rule adherence tendency and lower empathetic tendency.
@@ 5.1 Limitations and Future Directions
There are several limitations to our study. One limitation is that our method of simulating AI-Agents using only prompts designed from the BFI-2 falls short of capturing the intricate interplay of other important personality constructs, as well as individual backgrounds and societal complexity. We acknowledge that personality traits, while central to much of personality psychology, represent only one facet of what most people consider to be “personality.” Traits capture relatively stable patterns of thinking, feeling, and behaving, but they do not fully encompass other key components such as values, goals, roles, social identity, and emotional dynamics. Recent work has begun to investigate these other aspects in LLMs, such as how LLMs represent emotions [li2023large], human values [yao2025value], or identity [wang2025large]. These studies highlight that personality in artificial agents can, in principle, extend far beyond trait-level description. Building on these studies, a more integrative approach to LLM personality simulation could combine trait-based control with frameworks for modeling values, motivational systems, and identity-related processes. Such an approach would allow researchers to capture not only how an agent behaves across contexts, but also why it behaves that way, and how it might adapt in response to changing goals or social roles. Expanding in this direction could make AI-Agents a richer platform for studying complex aspects of human personality, bridging dispositional, motivational, and narrative analyses.
@@
Finally, due to the large number of conditions in our existing studies, we did not examine LLM parameters such as temperature and top-p; these were kept at their default values across our studies. Theoretically, these parameters could be optimized to generate responses that more closely align with human data. We are currently addressing this in a follow-up study on AI-Agents, using machine learning techniques to optimize these settings and improve response quality.